Blue Red Other
12 10 10
16 16 12
16 11 16
14 12 10
13 11 12
16 12 11

Given the random sample results above regarding favorite color and years of education, answer the following questions:

What is the within-groups degrees of freedom?
Answer #1

We have here, k = 3 and n = 3*6 = 18

The within-groups degrees of freedom = n-k = 18-3 = 15

i.e.  the within-groups degrees of freedom =15